OpenStudy (anonymous):

arithmetic sequence q : Determine a recursive de nition for each of the following sequence see equation... an=2n−3 (n=0,1,2...)

OpenStudy (anonymous):

\[a_1=-1\] \[a_2=1\] \[a_3=3\] \[a_4=5\]...

OpenStudy (anonymous):

sure looks like \[a_n=a_{n-1}+2\]

OpenStudy (anonymous):

that doesnt work for n=1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

this thing is just a line with slope 2. so of course you add 2 every time you go over 1.

OpenStudy (anonymous):

btw it does work for n = 1 yes? \[a_1=a_0+2\]

OpenStudy (anonymous):

is there a general way of working this out (for any sequence) i thought \[a_{0} = 0 so 0+2 ] is not -1

OpenStudy (anonymous):

i thought : \[a _{1} = 0 +2\] which isn't = -1

OpenStudy (anonymous):

Not sure how to write it like you guys but it looks like 2x-3

OpenStudy (anonymous):

\[a_n=2n-3\] Is that your formula?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so \[a_0=2\times 0-3=-3\]

OpenStudy (anonymous):

So, for instance you want 2nd term then you would plug in 2 for n 2(2)-3 4-3=1

OpenStudy (anonymous):

it is pretty clear that \[2n-3\] is a line with slope 2 yes? means every time n increases by 1, \[a_n\] increases by 2

OpenStudy (anonymous):

and y intercept of that line is at -3

OpenStudy (anonymous):

\[a_0=2\times 0-3=-3\] \[a_1=2\times 1-3=2-3=-1\] \[a_2=2\times 2-3=4-3=1\] \[a_3=2\times 3-3=6-3=3\] etc

OpenStudy (anonymous):

so as a recursion you could say \[a_0=-3\] \[a_n=a_{n-1}+2\]